OpenStudy (anonymous):

Erica solved the equation −5x − 25 = 78; her work is shown below. Identify the error and where it was made. −5x − 25 = 78 Step 1: −5x − 25 − 25 = 78 − 25 Step 2: −5x = 53 Step 3: negative five times x over negative five = fifty-three over negative five Step 4: x = negative fifty-three over five Step 1: She should have added 25 to each side. Step 1: She should have divided both sides by −5. Step 3: She should have divided both sides by −5. Step 3: She should have multiplied both sides by −5.

OpenStudy (anonymous):

Any idea?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i am not good at this

OpenStudy (anonymous):

are you still here

OpenStudy (anonymous):

Well when you're doing algebra the goal is to isolate a certain variable for these kinds of questions (meaning get x on its own) and to do this you usually do the opposite operation to move it to the other side, so for example if we have: 2x-3=45 our first step would be get move the -3 to the right side but what you do to one side you must do also do to the other. so add +3 on both sides and you can see it will get cancelled on the left 2x-3+3=45+3 2x=48 now you're left with 2x, you see "2x" is stuck together they are being multiplied, so we say ok what's the opposite operation that will remove the 2, oh right division! Divide both sides by 2 and you'll get x = 24

OpenStudy (anonymous):

Read that over, and see if you can figure it out, try your best and take your time and get back to me, it's ok if you get it wrong, or don't get a specific part, just ask.

OpenStudy (anonymous):

but what is the answer step one ore three please

OpenStudy (anonymous):

Well I want you to figure it out, I see you got rid of 2 answers which were wrong, so good job, can you get rid of 1 more?

OpenStudy (anonymous):

-5x-25=78 say this was a test, what would be your first step to isolate x, look at the example I gave you.

OpenStudy (anonymous):

is the first of step 3

OpenStudy (anonymous):

negative five times x over negative five = fifty-three over negative five , this?

OpenStudy (anonymous):

Are you sure James?

OpenStudy (anonymous):

nvm

OpenStudy (anonymous):

It's actually in my example if you just go over it

OpenStudy (anonymous):

the second of step 1

OpenStudy (anonymous):

I'm not sure what that means, are you saying step 1 is where the error is, where Erica was subtracting -25 instead of adding?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Yup, you're absolutely correct! :)

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

i need help with one more plz